OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ...

OpenStudy (anonymous):

Find a pattern. What are you doing to each number to produce the next in the sequence?

OpenStudy (anonymous):

plus 9

OpenStudy (anonymous):

-15 + 9(n - 1)

OpenStudy (danjs):

yeah, that type has constant change through one term to next

OpenStudy (anonymous):

so i'm right?

OpenStudy (danjs):

yeah you can test it, put in values of n, does it generate that sequence

OpenStudy (anonymous):

thank you!! can u help with another one

OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (danjs):

sure

OpenStudy (danjs):

that type of sequence, like i said has a cponstant change from one term to the next. like the +9 in the last one..

OpenStudy (anonymous):

\[a _{14}=-33 \] and \[a _{15} = 9\]

OpenStudy (danjs):

here they gave you consecutive terms, and told you the sequence is arithmetic

OpenStudy (danjs):

the constant change from one term to the next is 42

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[a _{n}\]= -579 + 42(n - 1) ??

OpenStudy (anonymous):

or is it an = -579 + 42(n + 1)

OpenStudy (danjs):

f the initial term is a1 and the common difference is d, then the nth term of the sequence is given by; an = a1 + (n-1)d - - - By taking the above result further, the nth term can be given also as; an = am + (n-m)d, where am is a random term in the sequence such that n > m.

OpenStudy (anonymous):

is my answer right then

OpenStudy (danjs):

an = am + (n-m)*d an = -33 + (n - 15) *42

OpenStudy (anonymous):

that isnt one of my answer choices

OpenStudy (danjs):

that more general form for any 2 terms, not just 1 and 2

OpenStudy (danjs):

yeah fix it up , an = initial term a0 + (n -1)*d

OpenStudy (danjs):

an = -33 + (n-14)*42 an = -33 + (n-1-13)*42 an = -33 + (n-1)*42 - (13*42) an = -579 + (n-1)*42

OpenStudy (danjs):

the initial term a0 is -579

OpenStudy (anonymous):

thx

OpenStudy (danjs):

welcome